OpenStudy (anonymous):

Divide: 25xy4 + 60xy3 ------------- 5xy

OpenStudy (anonymous):

5y3+12y2

OpenStudy (lgbasallote):

\[\large \frac{25xy^4 + 60xy^3}{5xy} \implies \frac{25xy^4}{5xy} + \frac{60xy^3}{5xy}\]

OpenStudy (anonymous):

That too